OpenStudy (anonymous):

in a right triangle find the elngth of the side not given b=2 c=|29

OpenStudy (anonymous):

a=5

OpenStudy (katielong):

that cant be logcally possible stevenator

OpenStudy (katielong):

29^2 - 2^2 = a^2 841 - 4 = 837 which is a^2 then find the square route of 837 which is 28.93095228

OpenStudy (katielong):

a = 28.93095228